Question 1. You will inject air into a sealed reaction bottle at room temperature and measure the pressure change within the bottle. Suppose 25.0 mL of air at an initial temperature of 20.5 degrees Celsius and an initial pressure of 1.010 atm was injected into a reaction bottle that contained 30.0 mL of water.

A. How many moles of air are contained in the syringe?

B. Pretend the gas you injected into the vessel all dissolved into the 30.0 mL of water. If this happened, what would the molar concentration of air be?

C. If the pressure in the flask is observed to be 1.175 atm with no change in temperature after the injection, what is the molar concentration-to-pressure change ratio? This value can be used to convert change in pressure within the vessel to a molar concentration of gas that escaped from the solution or vice versa.
Answer #1

Question 1.

You will inject air into a sealed reaction bottle at room temperature and measure the pressure change within the bottle.

Suppose 25.0 mL of air at an initial temperature of 20.5 degrees Celsius and an initial pressure of was injected into a reaction bottle that contained 30.0 mL of water.

A. How many moles of air are contained in the syringe?

using ideal gas law,

n = PV / RT = 1.010 atm * (25.0 mL* 1L/1000 mL) / 0.0821 L-atm/mol-K * 293.65 K = 0.00105 moles

B. Pretend the gas you injected into the vessel all dissolved into the 30.0 mL of water. If this happened, what would the molar concentration of air be?

molar concentration = moles/ Volume (in L) = 0.00105 mol / (30.0 mL* 1L/1000 mL) = 0.0350 mol/L

C. If the pressure in the flask is observed to be 1.175 atm with no change in temperature after the injection, what is the molar concentration-to-pressure change ratio?

molar concentration-to-pressure change ratio = 0.0350 M / ( 1.175 - 1) atm = 0.0350 M/ 0.175 atm = 0.2